Buses arrive randomly to a certain bus stop at an average rate of 10 per hour during the 10AM to 7PM time period. If you get to this bus stop at 1:00PM, what is the probability that the next bus arrives before 1:12PM

a).08647

b).1667

c).1353

d) no answer
Answer #1

a = 10 per hour = 10 per 60 min = 10/60 per 1 min = 1/6 per 1 min. Inter arrival time follows an exponential distribution wit